# Near-real-time flood inundation monitoring by Bayesian analysis for change point problems for Sentinel-1 time series

### v1

@@ Abstract
Keywords Synthetic Aperture Radar  $\cdot$ Flood monitoring  $\cdot$ Earth observation
@@ 1 Introduction
Flooding is one of the most frequent and devastating natural disasters globally [1, 2], resulting in substantial loss of life, damage to infrastructure, and severe economic consequences [3, 4, 5, 6]. The economic impact and human toll of flooding necessitate effective monitoring systems for early detection and rapid response [7, 8]. Near real-time flood inundation mapping, enabled by advances in remote sensing, has emerged as a critical tool for effective disaster response [7, 9, 10, 11]. These systems provide situational awareness to disaster responders, guide evacuation planning and search and rescue operations, and enable rapid damage assessments [12]. However, the widespread adoption of near real-time flood mapping remains limited due to challenges in remote sensing data analytics, creating an urgent need for more efficient and accessible monitoring approaches.
@@
Satellite remote sensing offers significant advantages for flood monitoring, including wide spatial coverage, high temporal resolution, and the ability to observe remote or inaccessible areas. However, optical remote sensing techniques face substantial limitations due to their inability to penetrate cloud cover, which is often present during flood events. Synthetic Aperture Radar (SAR) addresses these limitations by operating in the microwave portion of the electromagnetic spectrum, enabling all-weather, day-night imaging capability. The launch of the Sentinel-1 mission by the European Space Agency (ESA) has greatly enhanced the availability of SAR data, providing global coverage with a revisit time of 6-12 days and free data access. With high spatial resolution (5-20 m) and wide swath width (250 km), Sentinel-1 is well-suited for mapping flood extents at regional to continental scales [13].
@@
Despite SAR’s advantages for flood monitoring, significant challenges persist in developing fully automated, efficient, and accurate flood detection methods. Many existing approaches rely on supervised classification techniques that require extensive training data [14, 15, 16], which may not be readily available for all flood scenarios. Supervised classification methods require large amounts of labeled data to train algorithms, creating significant barriers to operational flood monitoring. The transferability of trained models to new flood scenarios or geographical regions is often limited [17], requiring additional data collection and model retraining. Furthermore, reliance on ancillary data (land cover maps, digital elevation models, hydrological information) can restrict the applicability and timeliness of flood mapping approaches, particularly in rapidly changing environments or regions with limited data infrastructure. Most existing time-series approaches for SAR-based flood monitoring rely either on simple change detection between pre- and post-flood images or on conventional statistical measures that don’t fully leverage the temporal information content in SAR data. Traditional change detection methods struggle to account for seasonal variations in backscatter or to distinguish between different types of change (e.g., flooding versus agricultural harvesting). While some Bayesian approaches have been applied in other remote sensing applications [18, 19], they have not been effectively adapted to exploit the unique characteristics of SAR time series for flood detection, particularly the need to identify abrupt changes against background variability.
@@
Thus, this study aims to overcome these limitations by developing an efficient, automated solution for near real-time flood monitoring using Sentinel-1 SAR time series data. We propose a pixel-wise Bayesian statistical model for SAR-based flood mapping that operates without the need for training data. Unlike existing approaches, our method leverages a Bayesian analysis of Change Point problem (BCP) as proposed by Barry and Hartigan [18] to identify flood-induced changes in the mean of our SAR time series data for the first time. Our workflow distinguishes between permanent water bodies and temporary flood inundations without requiring ancillary data by leveraging the temporal change characteristics in the probability domain rather than the raw backscatter domain. We use raw Sentinel-1 Ground Range Detected (GRD) time series as input without requiring training data, pre-trained models, or extensive manual intervention. The method achieves robust performance across diverse geographical contexts, making it particularly suitable for operational flood monitoring applications.
@@ 2.1 Flood monitoring by Optical remote sensing data
Optical satellite platforms such as Sentinel-2 and PlanetScope offer high spatial resolution and multispectral capabilities for flood mapping [20, 21]. However, cloud cover significantly impedes continuous monitoring during critical initial flood phases. Portalés-Julià et al. (2023) found approximately 50% of pixels obscured by clouds during the first post-flood satellite revisit [20]. Recent deep learning approaches have attempted to estimate water bodies beneath clouds [22, 14]. These models leverage spatial context from adjacent cloud-free pixels and temporal dependencies between acquisitions, achieving over 90% accuracy in water classification under 50-80% cloud coverage [20]. Despite these advances, challenges remain regarding training data requirements and geographical transferability.
@@ 2.2 Flood monitoring by remote sensing SAR data
SAR presents a compelling alternative due to its all-weather imaging capability [23, 24, 10]. SAR discriminates between water and non-water surfaces based on distinctive backscattering properties, with smooth water surfaces typically appearing dark due to specular reflection.
Interferometric SAR (InSAR) utilizes phase coherence information between acquisitions to detect inundation patterns, demonstrating effectiveness in complex urban environments [25]. However, the practical application of InSAR for timely flood monitoring is hindered by significant challenges. Normally, the extensive processing needed to derive these parameters often prevents real-time flood monitoring, limiting the method’s operational utility in urgent disaster response scenarios.
Deep learning methods have demonstrated exceptional capabilities for SAR-based flood detection [26, 27], with recent studies reporting classification accuracies exceeding 98% in certain contexts [16]. Specialized datasets like Sen1Flood11 [28] and Kuro-Siwo [29] have facilitated model training across diverse geographical contexts.
Nevertheless, significant challenges persist: (1) extensive labeled training data requirements, (2) limited model transferability across geographical contexts [15], (3) difficulty distinguishing between permanent water bodies and temporary inundation [30], and (4) complex backscatter interpretation in urban or vegetated areas [24].
@@ 2.3 Time Series Analysis for Flood Detection
Recent research has increasingly leveraged the temporal dimension of satellite data for improved flood detection [13]. Time series approaches exploit the sequential nature of acquisitions to identify anomalous patterns indicative of flood events, particularly effective with Sentinel-1’s consistent 6-12 day revisit cycle [7]. Schlaffer et al. (2015) proposed a harmonic analysis approach using multi-temporal SAR data to detect floods based on deviations from expected seasonal patterns [13]. Similarly, Tsyganskaya et al. (2018) and Tsyganskaya et al. (2019) employed time series analysis for wetland change detection and flooded vegetation mapping using Sentinel-1 data [31, 32]. These methods enable more robust discrimination between permanent water bodies and temporary inundation.
Conventional Bayesian methods offer advantages for hydrological applications by incorporating prior knowledge and quantifying uncertainty [18, 19]. D’Addabbo et al. (2016) developed a Bayesian network model fusing SAR and ancillary data [33], and Giordano et al. (2023) applied Bayesian statistical modeling to multi-temporal Sentinel-1 stacks for high-resolution flood monitoring [34]. However, existing Bayesian approaches primarily focus on retrospective analysis rather than near real-time detection. Most methods lack specific adaptations for the characteristic SAR backscatter patterns during flood events—sharp decreases during inundation followed by gradual increases during recession, limiting their effectiveness for timely flood monitoring.
@@ 3.1.1 UrbanSARFloods
As reference data, we utilized the UrbanSARFloods dataset, a comprehensive benchmark for flood detection using Sentinel-1 SAR data [17]. This dataset encompasses 8,879 non-overlapping image chips of 512×512 pixels at 20-m spatial resolution, covering an area of 807,500 km2 across 18 flood events on 5 continents. Every pixel in this dataset is classified into three distinct categories: non-flooded areas (0), flooded open areas (1), and flooded urban areas (2). UrbanSARFloods is systematically divided into training (4,501 chips), validation (1,970 chips), and testing (2,408 chips) sets, which are independent of flood events in training and validation data. This dataset ensures its quality by a two-tier annotation process, which combines semi-automatic labeling through conventional remote sensing techniques with manual verification using high-resolution optical data from PlanetScope and UAV imagery. While the BCP analysis does not require any training/validation set, we used them for the comparative analysis described in Section 4.
@@ 3.2 Bayesian analysis for change point problems
We implemented BCP using the bcp R package [19]. We selected the default tuning parameters of $\gamma=\lambda=0.2$ as recommended by Barry and Hartigan [18, 19]. Posterior means and probabilities were calculated based on 500 MCMC iterations after discarding the first 50 as burn-in, a configuration that achieved convergence while maintaining computational efficiency.
@@ 4.0.1 OTSU Thresholding Method
As a conventional baseline, we implemented the Otsu thresholding method [35]. This non-parametric technique determines an optimal threshold by maximizing the between-class variance between flood and non-flood pixels in SAR imagery. We applied the Otsu method to two different input configurations: VV polarization only and VH polarization only.
@@ 5.4 Quantitative Performance Evaluation
The table 3 summarizes the accuracy measures of our approach compared to OTSU thresholding and Siamese CNN approaches. For this evaluation, we used the globally optimazed parameters identified through sensitivity analysis (t=0.2, w=9) with VV-VH input. Our BCP approach achieved F1 scores ranging from 0.41 to 0.76, significantly outperforming OTSU thresholding and Siamese CNN approaches across all sites. The OTSU thresholding approach struggled with false positives, resulting in low precision across all sites. While the Siamese CNN showed improved performance over OTSU, it still fell short of our method’s capabilities.
@@ 6.1 Methodological Advances in SAR-Based Flood Detection
Our study introduces Bayesian analysis for change point problems (BCP) to near real-time flood monitoring applied to Sentinel-1 GRD time series data. Unlike supervised methods that require extensive training data and face challenges with geographical transferability [15, 16], our approach leverages the statistical properties of SAR time series to detect flood-induced changes without any training requirements. This fundamental training-free characteristic is a significant advantage, as it completely eliminates the overfitting issues inherent in deep learning and other machine learning approaches. Machine learning methods often perform well in areas represented in training datasets but fail when applied to novel geographical contexts or unusual flood patterns. By contrast, our BCP method requires no supervised learning process, making it immediately applicable to any region with sufficient SAR time series data, regardless of whether similar flood events have been previously observed and labeled.
Our method advances the field by reframing flood detection as a temporal changepoint identification problem rather than a binary classification task. Conventional classification approaches typically rely on absolute backscatter values or simple change metrics between two images, making them vulnerable to variations in instrument calibration, incidence angle effects, and seasonal changes in land cover [13]. In contrast, our Bayesian framework models the entire temporal trajectory of backscatter values, enabling more robust discrimination between normal fluctuations and flood-induced changes.
@@ 6.4 Limitations and Future Research Directions
The temporal resolution of Sentinel-1 (6-12 days) represents another limitation for monitoring rapidly evolving flood events [36]. Future research could explore the integration of our approach with other SAR platforms (e.g., RADARSAT Constellation Mission, ALOS-2) to achieve higher temporal frequency. Alternatively, data fusion techniques that combine SAR with optical or passive microwave observations could provide more continuous monitoring capabilities, leveraging the complementary strengths of different sensor types.
@@ 7 Conclusions
We present a novel approach to near real-time flood monitoring using Bayesian Changepoint Detection (BCP) with Sentinel-1 SAR time series data. Our key contributions include: (1) a training-free framework for automated flood detection, (2) innovative application of Bayesian analysis to SAR time series, and (3) empirical validation across diverse contexts. As climate change increases flood risks globally, our method contributes to disaster resilience and adaptation efforts through its accessibility, transferability, and operational efficiency. Our approach can be incorporated into semi-real-time flood monitoring systems due to its automatic processing capabilities. However, a key constraint is the infrequent coverage of Sentinel-1, which creates a bottleneck during urgent disaster response. Further research is needed to demonstrate our approach with higher-frequency SAR time series data.

### v2

@@ Abstract
keywords:
Synthetic Aperture Radar , Flood monitoring , Earth observation
journal: Remote Sensing of Environment
[…]
[not,tt,ns]organization=Graduate school of Science & Engineering, Saitama University, addressline=255 Shimo-Okubo, city=Sakura ward, Saitama, postcode=338-8570, state=Saitama, country=Japan
{graphicalabstract}[Uncaptioned image]
{highlights}
Novel Bayesian analysis for change point problems detects floods in Sentinel-1 time series
Method requires no training data or ancillary info to distinguish flood from permanent water
Automated processing of raw SAR data enables near real-time flood monitoring applications
@@ 1 Introduction
Flooding is one of the most frequent and devastating natural disasters globally [Alfieri2017-ma, Najibi2018-ta], resulting in substantial loss of life, damage to infrastructure, and severe economic consequences [Kundzewicz2014-ge, Jongman2021-ea, Tellman2021-hv, Jonkman2024-hi, Rentschler2021-dc]. The economic impact and human toll of flooding necessitate effective monitoring systems for early detection and rapid response [Oddo2019-or, Rentschler2020-pm]. Near real-time flood inundation mapping, enabled by advances in remote sensing, has emerged as a critical tool for effective disaster response [Oddo2019-or, Katiyar2021-fy, Shen2019-aj, Tay2020-cp]. These systems provide situational awareness to disaster responders, guide evacuation planning and search and rescue operations, and enable rapid damage assessments [Browder2021-io]. However, the widespread adoption of near real-time flood mapping remains limited due to challenges in remote sensing data analytics, creating an urgent need for more efficient and accessible monitoring approaches.
@@
Satellite remote sensing offers significant advantages for flood monitoring, including wide spatial coverage, high temporal resolution, and the ability to observe remote or inaccessible areas. However, optical remote sensing techniques face substantial limitations due to their inability to penetrate cloud cover, which is often present during flood events. Synthetic Aperture Radar (SAR) addresses these limitations by operating in the microwave portion of the electromagnetic spectrum, enabling all-weather, day-night imaging capability. The launch of the Sentinel-1 mission by the European Space Agency (ESA) has greatly enhanced the availability of SAR data, providing global coverage with a revisit time of 6-12 days and free data access. With high spatial resolution (5-20 m) and wide swath width (250 km), Sentinel-1 is well-suited for mapping flood extents at regional to continental scales [Schlaffer2015-uv].
@@
Despite SAR’s advantages for flood monitoring, significant challenges persist in developing fully automated, efficient, and accurate flood detection methods. Many existing approaches rely on supervised classification techniques that require extensive training data [Shastry2023-wl, Wu2023-kc, Ghosh2024-ub], which may not be readily available for all flood scenarios. Supervised classification methods require large amounts of labeled data to train algorithms, creating significant barriers to operational flood monitoring. The transferability of trained models to new flood scenarios or geographical regions is often limited [Zhao2024-lv], requiring additional data collection and model retraining. Furthermore, reliance on ancillary data (land cover maps, digital elevation models, hydrological information) can restrict the applicability and timeliness of flood mapping approaches, particularly in rapidly changing environments or regions with limited data infrastructure. Most existing time-series approaches for SAR-based flood monitoring rely either on simple change detection between pre- and post-flood images or on conventional statistical measures that don’t fully leverage the temporal information content in SAR data. Traditional change detection methods struggle to account for seasonal variations in backscatter or to distinguish between different types of change (e.g., flooding versus agricultural harvesting). While some Bayesian approaches have been applied in other remote sensing applications [Barry1993-qp, Erdman2008-kv], they have not been effectively adapted to exploit the unique characteristics of SAR time series for flood detection, particularly the need to identify abrupt changes against background variability.
@@
Thus, this study aims to overcome these limitations by developing an efficient, automated solution for near real-time flood monitoring using Sentinel-1 SAR time series data. We propose a pixel-wise Bayesian statistical model for SAR-based flood mapping that operates without the need for training data. Unlike existing approaches, our method leverages a Bayesian analysis of Change Point problem (BCP) as proposed by Barry and Hartigan [Barry1993-qp] to identify flood-induced changes in the mean of our SAR time series data for the first time. Our workflow distinguishes between permanent water bodies and temporary flood inundations without requiring ancillary data by leveraging the temporal change characteristics in the probability domain rather than the raw backscatter domain. We use raw Sentinel-1 Ground Range Detected (GRD) time series as input without requiring training data, pre-trained models, or extensive manual intervention. The method achieves robust performance across diverse geographical contexts, making it particularly suitable for operational flood monitoring applications.
@@ 2.1 Flood monitoring by Optical remote sensing data
Optical satellite platforms such as Sentinel-2 and PlanetScope offer high spatial resolution and multispectral capabilities for flood mapping [Portales-Julia2023-pz, Wieland2019-je]. However, cloud cover significantly impedes continuous monitoring during critical initial flood phases. Portalés-Julià et al. (2023) found approximately 50% of pixels obscured by clouds during the first post-flood satellite revisit [Portales-Julia2023-pz]. Recent deep learning approaches have attempted to estimate water bodies beneath clouds [Li2021-fq, Shastry2023-wl]. These models leverage spatial context from adjacent cloud-free pixels and temporal dependencies between acquisitions, achieving over 90% accuracy in water classification under 50-80% cloud coverage [Portales-Julia2023-pz]. Despite these advances, challenges remain regarding training data requirements and geographical transferability.
@@ 2.2 Flood monitoring by remote sensing SAR data
SAR presents a compelling alternative due to its all-weather imaging capability [Landuyt2019-cz, Amitrano2024-av, Shen2019-aj]. SAR discriminates between water and non-water surfaces based on distinctive backscattering properties, with smooth water surfaces typically appearing dark due to specular reflection.
Interferometric SAR (InSAR) utilizes phase coherence information between acquisitions to detect inundation patterns, demonstrating effectiveness in complex urban environments [Chini2019-me]. However, the practical application of InSAR for timely flood monitoring is hindered by significant challenges. Normally, the extensive processing needed to derive these parameters often prevents real-time flood monitoring, limiting the method’s operational utility in urgent disaster response scenarios.
Deep learning methods have demonstrated exceptional capabilities for SAR-based flood detection [Saleh2024-vd, Bereczky2022-qy], with recent studies reporting classification accuracies exceeding 98% in certain contexts [Ghosh2024-ub]. Specialized datasets like Sen1Flood11 [Bonafilia2020-sk] and Kuro-Siwo [Bountos2023-jg] have facilitated model training across diverse geographical contexts.
Nevertheless, significant challenges persist: (1) extensive labeled training data requirements, (2) limited model transferability across geographical contexts [Wu2023-kc], (3) difficulty distinguishing between permanent water bodies and temporary inundation [Notti2018-vs], and (4) complex backscatter interpretation in urban or vegetated areas [Amitrano2024-av].
@@ 2.3 Time Series Analysis for Flood Detection
Recent research has increasingly leveraged the temporal dimension of satellite data for improved flood detection [Schlaffer2015-uv]. Time series approaches exploit the sequential nature of acquisitions to identify anomalous patterns indicative of flood events, particularly effective with Sentinel-1’s consistent 6-12 day revisit cycle [Oddo2019-or]. Schlaffer et al. (2015) proposed a harmonic analysis approach using multi-temporal SAR data to detect floods based on deviations from expected seasonal patterns [Schlaffer2015-uv]. Similarly, Tsyganskaya et al. (2018) and Tsyganskaya et al. (2019) employed time series analysis for wetland change detection and flooded vegetation mapping using Sentinel-1 data [Tsyganskaya2018-ts, Tsyganskaya2019-nv]. These methods enable more robust discrimination between permanent water bodies and temporary inundation.
Conventional Bayesian methods offer advantages for hydrological applications by incorporating prior knowledge and quantifying uncertainty [Barry1993-qp, Erdman2008-kv]. D’Addabbo et al. (2016) developed a Bayesian network model fusing SAR and ancillary data [D-Addabbo2016-yi], and Giordano et al. (2023) applied Bayesian statistical modeling to multi-temporal Sentinel-1 stacks for high-resolution flood monitoring [Giordano2023-ys]. However, existing Bayesian approaches primarily focus on retrospective analysis rather than near real-time detection. Most methods lack specific adaptations for the characteristic SAR backscatter patterns during flood events—sharp decreases during inundation followed by gradual increases during recession, limiting their effectiveness for timely flood monitoring.
@@ 3.1.1 UrbanSARFloods
As reference data, we utilized the UrbanSARFloods dataset, a comprehensive benchmark for flood detection using Sentinel-1 SAR data [Zhao2024-lv]. This dataset encompasses 8,879 non-overlapping image chips of 512×512 pixels at 20-m spatial resolution, covering an area of 807,500 km2 across 18 flood events on 5 continents. Every pixel in this dataset is classified into three distinct categories: non-flooded areas (0), flooded open areas (1), and flooded urban areas (2). UrbanSARFloods is systematically divided into training (4,501 chips), validation (1,970 chips), and testing (2,408 chips) sets, which are independent of flood events in training and validation data. This dataset ensures its quality by a two-tier annotation process, which combines semi-automatic labeling through conventional remote sensing techniques with manual verification using high-resolution optical data from PlanetScope and UAV imagery. While the BCP analysis does not require any training/validation set, we used them for the comparative analysis described in Section 4.
@@ 3.2 Bayesian analysis for change point problems
We implemented BCP using the bcp R package [Erdman2008-kv]. We selected the default tuning parameters of $\gamma=\lambda=0.2$ as recommended by Barry and Hartigan [Barry1993-qp, Erdman2008-kv]. Posterior means and probabilities were calculated based on 500 MCMC iterations after discarding the first 50 as burn-in, a configuration that achieved convergence while maintaining computational efficiency.
@@ 4.0.1 OTSU Thresholding Method
As a conventional baseline, we implemented the Otsu thresholding method [Otsu1979-dj]. This non-parametric technique determines an optimal threshold by maximizing the between-class variance between flood and non-flood pixels in SAR imagery. We applied the Otsu method to two different input configurations: VV polarization only and VH polarization only.
@@ 5.4 Quantitative Performance Evaluation
The table LABEL:tab:comparison summarizes the accuracy measures of our approach compared to OTSU thresholding and Siamese CNN approaches. For this evaluation, we used the globally optimazed parameters identified through sensitivity analysis (t=0.2, w=9) with VV-VH input. Our BCP approach achieved F1 scores ranging from 0.41 to 0.76, significantly outperforming OTSU thresholding and Siamese CNN approaches across all sites. The OTSU thresholding approach struggled with false positives, resulting in low precision across all sites. While the Siamese CNN showed improved performance over OTSU, it still fell short of our method’s capabilities.
@@ 6.1 Methodological Advances in SAR-Based Flood Detection
Our study introduces Bayesian analysis for change point problems (BCP) to near real-time flood monitoring applied to Sentinel-1 GRD time series data. Unlike supervised methods that require extensive training data and face challenges with geographical transferability [Wu2023-kc, Ghosh2024-ub], our approach leverages the statistical properties of SAR time series to detect flood-induced changes without any training requirements. This fundamental training-free characteristic is a significant advantage, as it completely eliminates the overfitting issues inherent in deep learning and other machine learning approaches. Machine learning methods often perform well in areas represented in training datasets but fail when applied to novel geographical contexts or unusual flood patterns. By contrast, our BCP method requires no supervised learning process, making it immediately applicable to any region with sufficient SAR time series data, regardless of whether similar flood events have been previously observed and labeled.
Our method advances the field by reframing flood detection as a temporal changepoint identification problem rather than a binary classification task. Conventional classification approaches typically rely on absolute backscatter values or simple change metrics between two images, making them vulnerable to variations in instrument calibration, incidence angle effects, and seasonal changes in land cover [Schlaffer2015-uv]. In contrast, our Bayesian framework models the entire temporal trajectory of backscatter values, enabling more robust discrimination between normal fluctuations and flood-induced changes.
@@ 6.4 Limitations and Future Research Directions
The temporal resolution of Sentinel-1 (6-12 days) represents another limitation for monitoring rapidly evolving flood events [Munasinghe2023-oc]. Future research could explore the integration of our approach with other SAR platforms (e.g., RADARSAT Constellation Mission, ALOS-2) to achieve higher temporal frequency. Alternatively, data fusion techniques that combine SAR with optical or passive microwave observations could provide more continuous monitoring capabilities, leveraging the complementary strengths of different sensor types.
@@ 7 Conclusions
We present a novel approach to near real-time flood monitoring using Bayesian Changepoint Detection (BCP) with Sentinel-1 SAR time series data. Our key contributions include: (1) a training-free framework for automated flood detection, (2) innovative application of Bayesian analysis to SAR time series, and (3) empirical validation across diverse contexts.
As climate change increases flood risks globally, our method contributes to disaster resilience and adaptation efforts through its accessibility, transferability, and operational efficiency.
Our approach can be incorporated into semi-real-time flood monitoring systems due to its automatic processing capabilities. However, a key constraint is the infrequent coverage of Sentinel-1, which creates a bottleneck during urgent disaster response. Further research is needed to demonstrate our approach with higher-frequency SAR time series data.